# GRAMMAR: Grounded and Modular Methodology for Assessment of Domain-Specific Retrieval-Augmented Language Models

### v4

@@ Abstract
Retrieval-augmented Generation (RAG) systems have been actively studied and deployed across various industries to query on domain-specific knowledge base. However, evaluating these systems presents unique challenges due to the scarcity of domain-specific queries and corresponding ground truths, as well as a lack of systematic approaches to diagnosing the cause of failure cases—whether they stem from knowledge deficits or issues related to system robustness. To address these challenges, we introduce GRAMMAR (GRounded And Modular Methodology for Assessment of RAG), an evaluation framework comprising two key elements: 1) a data generation process that leverages relational databases and LLMs to efficiently produce scalable query-answer pairs. This method facilitates the separation of query logic from linguistic variations for enhanced debugging capabilities; and 2) an evaluation framework that differentiates knowledge gaps from robustness and enables the identification of defective modules. Our empirical results underscore the limitations of current reference-free evaluation approaches and the reliability of GRAMMAR to accurately identify model vulnerabilities.
GRAMMAR: Grounded and Modular Methodology for Assessment of Domain-Specific Retrieval-Augmented Language Models
Figure 1: The dual-phase evaluation approach for the GRAMMAR model. The left section outlines the data generation process for creating sets of hypothetically robust ($\mathcal{D}_{\text{robust}}$) and non-robust ($\mathcal{D}_{\text{non-robust}}$) data to test model hypotheses. The right section details the methodology for identifying defective modules. This grouping step accumulates queries with identical semantics and delineates three group types for identifying detective modules. Gap Groups indicate a complete lack of correct responses, reflecting deficiencies in the model’s knowledge or capabilities. Robust Groups show the model’s comprehensive understanding, with correct responses to all queries in the subset. Non-robust Groups highlight the robustness issue, where the model correctly answers some but not all text forms with identical query logic.
@@ 1 Introduction
The emergent ability of Large Language Models (LLMs) has advanced a active research and deployment of Retrieval-Augmented Generation (RAG) systems for domain-specific question answering. This trend underscores the critical need for reliable evaluation of these systems, a task that remains intricate and largely unaddressed.
@@
One of the foremost challenges is the scarcity of domain-specific data. Typically, evaluators resort to using data designed for human assessment, as referenced in studies such as Santurkar et al. (2023); Wang et al. (2022); Zhong et al. (2022, 2023); Hendrycks et al. (2021); Choi et al. (2023), which mainly focuses on evaluating commonsense and broad world knowledge. However, the assessment of industrial RAG systems demands knowledge specific to particular domains, such as details about company projects and employees. While user queries can be gathered at the time of system deployment, acquiring accurate ground truths poses a significant challenge. Recent studies have explored the use of reference-free LLM evaluators in the absence of ground truths (Chern et al., 2023; Min et al., 2023; Es et al., 2023). Nonetheless, the reliability of such reference-free evaluation methods remains questionable. Our analysis of two representative methods convincingly demonstrates their false assessment across accurate and inaccurate predictions: only RAGAS demonstrates reliability in identifying correct predictions. In this paper, an automatic data generation methodology is proposed to ensure the accessibility of ground truths (Contribution 1). The methodology leverages relational databases and LLMs to enable the extraction of ground truths via SQL queries.
A second significant challenge relates to the robustness of models (Contribution 2). Specifically, we observe inconsistencies in the models’ responses to queries that, while semantically equivalent, differ in their linguistic presentation, as highlighted in Shen et al. (2023). Such inconsistencies pose a direct challenge to the crucial requirement for consistent performance. Researchers often hypothesize about the factors influencing the robustness or susceptibility of retrieval and language models to certain attributes, questioning what impedes these models from responding accurately to queries. However, the current body of research, including works on adversarial robustness (Alzantot et al., 2018; Li et al., 2020, 2019), lacks the mechanisms to effectively test these hypotheses. To address this gap, we introduce a evaluation workflow, as domonstrated in Figure 1: 1) utilizing the proposed data generation process that creates hypothetically robust data $\mathcal{D}_{\text{robust}}$ and non-robust data $\mathcal{D}_{\text{non-robust}}$ according to a proposed hypothesis, 2) identifying defective modules based on a grouping mechanism, contextual sufficiency analysis and a reliable robustness metric that minimizes knowledge gap interference. Our empirical study highlights the reliability of GRAMMAR to verify the hypothesis and further identify the defective modules.
@@ 2 Background
Retrieval-augmented Generation
@@
RAG systems augment LLMs with retrieval for domain-specific use where extensive databases can potentially provide necessary information. A retrieval can be used for retrieving relevant passages or documents. A dense retrieval (Mialon et al., 2023; Lewis et al., 2020; Borgeaud et al., 2022) consists of an embedding model to produce a vector $v_{q}$ for any given query $q$. This vector is then used to identify pertinent text segments normally via dot product calculations with the vectors of text chunks. The most relevant chunks are aggregated to form an additional context $c$, which serves as the factual basis for the response generation, until adding another chunk would surpass the maximum allowable context length. Subsequently, an LLM will process this aggregated context alongside the original query to generate a predicted answer $\hat{a}$ for $q$. In contrast, sparse retrieval computes similarity without parametric embedding models. Keyword matching . For example, sparse retrieval techniques operate without the need for parametric embedding models. For instance, keyword-matching approaches (Manning, 2009) select documents that share the highest number of common words for context retrieval. Additionally, methods like TF-IDF (Chen et al., 2017) factor in the inverse document frequency, emphasizing the significance of less common words which are likely to be more indicative of pertinent content.
Knowledge-intensive Evaluation
Many evaluation protocols of RAG and LLMs focus on open-domain knowledge. Both the queries $Q$ and their ground-truth answers $A$ in the open domains can be easily accessed via online forums, public surveys (Santurkar et al., 2023) and qualification and admission exams that were originally designed for humans Wang et al. (2022); Zhong et al. (2022, 2023); Hendrycks et al. (2021); Choi et al. (2023). Since LLMs are originally trained on open-domain knowledge, these evaluation protocols on RAG focuses more on whether LLMs’ factuality can be improved by retrieval. However, this paper focuses on industrial scenarios where retrieval mostly provides new knowledge to LLMs and ground-truth answers are hardly available. The reference-free evaluation protocol is commonly proposed for this situation, e.g., ARES (Saad-Falcon et al., 2023) and RAGAS (Es et al., 2023). For example, instead of the correctness of answers against the users’ queries, Saad-Falcon et al. (2023) assess answer faithfulness against the context from retrieval modules by profiling LLMs to generate evaluation scores. The results can be skeptical due to the reliability of retrieval and the fairness of LLMs as evaluators. Reference-free evaluation is also utilized to assess LLMs’ open-domain knowledge (Kadavath et al., 2022). SelfCheck (Manakul et al., 2023) relies on the stochastic nature of LLMs for validations, while FactScore (Min et al., 2023) and FacTool (Chern et al., 2023) use Retrieval+LLM as a validation system. FactScore and FactTool are naturally not suitable for evaluating RAG systems that already include retrieval as an inherent component.
3 GRAMMAR: Grounded and Modular Methodology for Assessment of RAG
GRAMMAR is structured around two key components: 1) Grounded Data Generation Method: Grounded Data Generation Method: This approach generates text queries and corresponding answers from an existing knowledge base (relational database) for targeted evaluations. Using SQL as an intermediate representation, it allows for the controlled creation of various text queries with consistent underlying logic. The process and its implementation are detailed in § 3.1 and 3.2, and illustrated in Figure 2. 2) Modular Evaluation Protocol: This protocol assesses robustness by not only identifying proposed vulnerabilities but also pinpointing which component exhibits robustness issues against these vulnerabilities, as detailed in § 3.3.
Note that the modular evaluation relies on the data generation method. The data generation method produces contrastive data clusters with the same query logic, ensuring a fair evaluation (See Figure 1). Another key aspect of the design choice is the significant combinatorial potential inherent in the proposed data generation process, influenced by various factors. Specifically, a manually crafted database with a modest number of attributes and rows suffices for testing purposes, given the extensive variety of query permutations made possible through the combinatorial approach. A comprehensive analysis of these concepts is presented in Appendix B.3 and C.
Algorithm 1 Grounded Data Generation Process.
@@ 3 GRAMMAR: Grounded and Modular Methodology for Assessment of RAG
1: SQL Template Generator $g_{\mathrm{sql}}$, Text Template Generator $g_{\mathrm{t}}$, Semantic criteria for SQL template generation $\mathcal{C}_{\mathrm{sql}}$, Linguistic criteria for text template generation $\mathcal{C}_{\mathrm{t}}$, Database $\mathcal{D}$, Database Schema $\mathcal{S}=\left\{S_{1},S_{2},\ldots,S_{n}\right\}$, where each $S_{i}$ is a schema for a table
2:$Q\leftarrow\emptyset$triangleright Used to store final evaluation data
3:$\mathcal{S}_{\text{target}}\in\mathcal{S}$triangleright Step 1: Define schema of a domain (§3.1)
4:$\{tpl_{\text{sql}}\}\leftarrow g_{\mathrm{sql}}(\mathcal{S}_{\text{target}},\mathcal{C}_{\mathrm{sql}})$triangleright Step 2: Generate SQL templates applying semantic criteria (§3.1)
5:for $tpl_{\mathrm{sql}}\in\{tpl_{\text{sql}}\}$ do
6:     $\{tpl_{\text{t}}\}\leftarrow g_{\mathrm{t}}(tpl_{\mathrm{sql}},\mathcal{C}_{\mathrm{t}})$triangleright Step 3: Generate text templates applying linguistic criteria (§3.1))
7:     $P\leftarrow$ extract placeholders from $tpl_{\mathrm{sql}}$
8:     for each $p\in P$ do
9:         $C\leftarrow$ extract column from $p$
10:         $V_{p}\leftarrow$ query $\mathcal{D}$ for distinct values of $C$
11:     end for
12:     $Comb\leftarrow$ Cartesian product of $\{V_{p}:p\in P\}$
13:     for each $comb\in Comb$ do
14:         $q_{\mathrm{sql}}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{sql}}$triangleright Step 4: Generate SQL queries (§3.2)
15:         $a\leftarrow$ query $\mathcal{D}(q_{\mathrm{sql}})$triangleright Step 5: Query DB for answers (§3.2)
16:         $\{q_{t}\}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{t}}\in\{tpl_{\text{t}}\}$triangleright Step 6: List of text queries (§3.2)
17:         $Q\leftarrow Q\cup\{(\{q_{t}\},a)\}$
@@ 3.1 Generating Query Templates
This process begins by taking a database schema as input and subsequently produces SQL query templates along with their corresponding textual representations.
@@
Database schema serves as the blueprint that defines how classes of entities are organized and the relations among them. Appendix A demonstrates the reasons why we use database schema for knowledge representation. Let a database schema with a collection of $N$ tables $S=\left\{S_{1},S_{2},\ldots,S_{N}\right\}$, where each $S_{i}$ is a schema for a table. A table schema $S_{i}$ consists of the table name $T_{i}$, a set of attributes $\left\{A_{1},A_{2},\ldots,A_{M}\right\}$ and a set of constraints, i.e., $S_{i}=\left(T_{i},\left\{A_{1},A_{2},\ldots,A_{m}\right\}\right.,C)$, where the constraint $C$ demonstrates the primary key $PK$ and foreign keys $FK$. Specifically, $FK\left(T_{i}.A_{k}\rightarrow T_{j}.A_{h}\right)$ indicates a foreign key $A_{k}$ in table $T_{i}$ referencing the primary key $A_{h}$ in table $T_{j}$. The schema defines the knowledge structure of an entity, which can be enough to generate query logic that is not tied to any specific linguistic expressions and sensitive data.
@@
$tpl_{\mathrm{sql}}$ can be instantiated into various text templates, each denoted as $tpl_{t}$. This process, driven by the text template generator $g_{\mathrm{t}}$, leverages linguistic criteria $\mathcal{C}_{\mathrm{t}}$ to transform $tpl_{\mathrm{sql}}$ into natural language forms that align with specified linguistic characteristics, e.g., complexity, length and stylistic nuances. The text template generator $g_{\mathrm{t}}$ is also operationalized using a sophisticated language model, such as GPT-4, which is adept at producing diverse linguistic variations of the same semantic content.
$\{tpl_{\text{t}}\}=g_{\text{t}}(tpl_{\text{sql}},\mathcal{C}_{\text{t}})$
, where $\{tpl_{\text{t}}\}$ signifies the resulting list of text templates.

### v6

# GRAMMAR: Grounded and Modular Methodology for Assessment of Closed-Domain Retrieval-Augmented Language Models

@@ Abstract
Retrieval-Augmented Generation (RAG) systems are widely used across various industries for querying closed-domain and in-house knowledge bases. However, evaluating these systems presents significant challenges due to the private nature of closed-domain data and a scarcity of queries with verifiable ground truths. Moreover, there is a lack of analytical methods to diagnose problematic modules and identify types of failure, such as those caused by knowledge deficits or issues with robustness. To address these challenges, we introduce GRAMMAR (GRounded And Modular Methodology for Assessment of RAG), an evaluation framework comprising a grounded data generation process and an evaluation protocol that effectively pinpoints defective modules. Our validation experiments reveal that GRAMMAR provides a reliable approach for identifying vulnerable modules and supports hypothesis testing for textual form vulnerabilities. An open-source tool accompanying this framework is available in our GitHub repository https://github.com/xinzhel/grammar, allowing for easy reproduction of our results and enabling reliable and modular evaluation in closed-domain settings.
GRAMMAR: Grounded and Modular Methodology for Assessment of Closed-Domain Retrieval-Augmented Language Models
[…]
@@ 1 Introduction
The emergent capabilities of Large Language Models (LLMs) have driven significant research and the widespread deployment of Retrieval-Augmented Generation (RAG) systems for closed-domain settings. This growing trend highlights the critical need for reliable evaluation methods tailored to RAG systems.
@@
A key challenge in evaluating RAG systems is the collection of domain-specific data that includes accurate ground truths. Researchers often rely on data originally designed for human assessment, as seen in studies such as Santurkar et al. (2023); Wang et al. (2022); Zhong et al. (2022, 2023); Hendrycks et al. (2021); Choi et al. (2023). However, these data sources typically reflect open-domain commonsense and world knowledge, whereas industrial in-house RAG systems require data that captures closed-domain knowledge, such as details about company projects and employees. Although user queries can be gathered during system deployment, obtaining accurate ground truths remains a significant challenge. To address this, recent studies have explored reference-free LLM evaluators for cases where ground truths are unavailable (Chern et al., 2023; Min et al., 2023; Es et al., 2023). However, the reliability of these reference-free evaluation methods is still questionable.
Another critical challenge is assessing robustness in RAG systems, where a system may possess adequate knowledge for the semantics of a query but respond inconsistently to different query forms, as highlighted in Shen et al. (2023). Developers and researchers often hypothesize about factors that could impact robustness, such as vulnerable modules (e.g., retrieval mechanisms or language models) or specific input attributes. However, existing research, e.g., studies on adversarial robustness (Alzantot et al., 2018; Li et al., 2020, 2019), lacks effective tools for identifying non-robust modules and analytically testing these hypotheses.
Recognizing these challenges, the paper makes the following contributions:
Proposing a grounded, controllable data generation process (GRAMMAR-Gen). To address the issue of ground truth scarcity, we propose GRAMMAR-Gen, a grounded data-generation process designed to ensure reliable evaluation. This process leverages relational databases and LLMs to extract ground truths via SQL queries. Additionally, the template-based query generation enables scalable data creation. By leveraging the capabilities of LLMs, it allows the controlled generation of data with diverse linguistic attributes. Notably, a similar data-generation process for evaluating domain knowledge was proposed concurrently by Tu et al. (2024) [^1] It was published after the initial preprint/submission of this work..
Proposing a protocol for modular evaluation and hypothesis testing (GRAMMAR-Eval). GRAMMAR-Eval combines a grouping and tagging mechanism based on GRAMMAR-Gen with effective strategies to both identify knowledge gaps in the retrieval database and determine whether robustness issues stem from the retrieval mechanism or the language models (LMs). Our empirical results demonstrate the effectiveness of GRAMMAR-Eval in identifying deliberately bugged retrieval systems and accurately testing hypotheses about which input attributes contribute to robustness issues. An example of this process is illustrated in Figure 1.
@@ 2 Background
Evaluating the performance of Retrieval-Augmented Generation (RAG) systems, especially in closed-domain settings, poses unique challenges, leading to an absence of reliable reference-based evaluation protocols. This section outlines existing approaches and highlights the gaps addressed by our proposed framework, GRAMMAR.
Retrieval-augmented Generation (RAG)
@@
RAG systems augment LLMs with retrieval for domain-specific use where extensive databases can potentially provide necessary information. A retrieval can be used for retrieving relevant passages or documents. A dense retrieval (Mialon et al., 2023; Lewis et al., 2020; Borgeaud et al., 2022) consists of an embedding model to produce a vector $v_{q}$ for any given query $q$. This vector $v_{q}$ is then used to identify pertinent text segments normally via dot product calculations with the vectors of text chunks. The most relevant chunks are aggregated to form an additional context $c$, which serves as the factual basis for the response generation, until adding another chunk would surpass the maximum allowable context length. Subsequently, an LLM will process this aggregated context alongside the original query $q$ to generate a predicted answer $\hat{a}$ for it.
In contrast, sparse retrieval computes similarity without parametric embedding models. For instance, keyword-matching approaches (Manning, 2009) select documents that share the highest number of common words for context retrieval. Additionally, methods like TF-IDF (Chen et al., 2017) factor in the inverse document frequency, emphasizing the significance of less common words which are likely to be more indicative of pertinent content.
Reference-Free Evaluation
Reference-free evaluation methods (Es et al., 2023) assess model performance without requiring predefined ground truths. These methods can be categorized according to their applicable systems: 1) Protocols for LLMs: FactScore (Min et al., 2023) and FacTool (Chern et al., 2023) use retrieval combined with LLMs as validators, but these are not directly applicable to RAG systems as retrieval is reserved for evaluation. 2) Protocols for RAG: ARES and RAGAS focus on evaluating the faithfulness of generated answers against retrieved contexts, with LLMs scoring the relevance and correctness of the context (Saad-Falcon et al., 2023; Es et al., 2023). 3) Protocols for Both: SelfCheck (Manakul et al., 2023) utilizes the stochastic nature of LLMs for self-validation. While these methods offer some insights, their reliability is often questioned, especially for closed-domain applications where specific, accurate answers are crucial.
Reference-Based Evaluation
While reference-free evaluation can be applied for evaluating both open-domain and closed-domain queries, reference-based evaluation typically assesses open-domain knowledge. The reason is that open-domain queries are widely available with ground-truth answers from public resources, exams, and surveys (Santurkar et al., 2023; Wang et al., 2022; Zhong et al., 2022, 2023; Hendrycks et al., 2021; Choi et al., 2023). These evaluations compare generated answers to known correct answers, calculating metrics that reflect the model’s accuracy. However, closed-domain scenarios lack reference-based evaluation due to the unavailability of ground-truth answers. GRAMMAR-Gen is proposed to solve this issue.
3 Is Reference-free Evaluation Reliable?
This section analyzes the reliability of reference-free evaluation methods using 198 examples from an engineering company. The RAG system for evaluation is implemented with dense retrieval and GPT-3.5  [^2] An OpenAI embedding model for document embedding and gpt-3.5-turbo-16k.
This preliminary study briefly highlights the potential benefits of reference-based evaluation, as our primary contribution focuses on proposing a reference-based evaluation framework. A detailed analysis of reference-free evaluation on state-of-the-art models is left for future work.
Two Evaluation Perspectives: Optimism and Cynicism
To structure this analysis, reference-free evaluation is treated as a binary classification task. The ground truth reflects the actual correctness of RAG responses, while the evaluation models’ judgments serve as the task’s predictions. Thus, low precision indicates an optimistic bias (where incorrect RAG responses are often judged as correct), and low recall indicates a cynical bias (where correct RAG responses are often judged as incorrect).
Two Evaluation Protocols
We assess two reference-free protocols.
  • RAGAS-Fact (Es et al., 2023): This protocol utilizes the context-query-response triplets to assess the veracity of responses. It evaluates the faithfulness of a response by calculating the ratio of claims grounded on the context to the total claims made. This process involves identifying statements that hold atomic facts, following the methodology outlined by Chern et al. (2023).
  • SelfCheck (Manakul et al., 2023): This protocol relies on the stochastic generation of responses, based on the premise that incorrect answers are unlikely to be consistently produced. This principle, initially applied to LLMs, is adapted in our analysis of RAG systems. In a departure from its original application, we enhance the evaluation prompt to include not only stochastic responses but also the query itself. Refer to Appendix A for details. To generate four stochastic samples, we adjust the temperature setting to 1.0, contrasting with a temperature of 0.0 used to generate the primary response. A response is deemed correct if it aligns consistently across all stochastic samples.
Results: Both are Extremely Optimistic on Wrong Predictions While SelfCheck Becomes Too Cynical on Correct Predictions
As shown in Table 1, both RAGAS and SelfCheck achieve low precision (19% and 15%, respectively), highlighting a deficiency in correctly identifying erroneous predictions. For RAGAS, this may be because the RAG system produces responses that, while contextually relevant, fail to directly address the intended query. However, RAGAS demonstrates high recall, indicating accurate assessment of correct RAG responses. Overall, the preliminary insight suggests that reference-free evaluation methods are not reliable for evaluation, underscoring the critical need for a robust, reference-based evaluation framework to ensure reliable model assessment.
Precision Recall
RAGAS-Fact 19% (11-27%) 92% (86-97%)
SelfCheck 15% (0.08-0.22%) 50% (0.4-0.59%)
Table 1: Reliability of reference-free evaluation protocols. Taking the size of the sample into account, the Z-test with 95% confidence intervals is utilized.
4 GRAMMAR
This section introduces GRAMMAR-Gen and GRAMMAR-Eval for grounded and modular evaluation: 1) GRAMMAR-Gen begins by generating templates based on a database schema, enabling scalable data creation  [^3] Further details can be found in Appendices D.2 and E. By leveraging LLMs, GRAMMAR-Gen offers controllability in generating text variations, which not only supports scalable data generation but also facilitates modular robustness evaluation and hypothesis testing. The process of SQL and text template generation is detailed in § 4.1. SQL queries, serving as intermediate representations, are used to create ground-truth answers by querying relational databases, as explained in § 4.2. The overall process is illustrated in Algorithm 1 and Figure 2. 2) GRAMMAR-Eval, an evaluation protocol that effectively pinpoints defective modules by assessing the knowledge deficit of the retrieval database and detecting robustness issues originating from either the retrieval mechanisms or the language models. Further details can be found in § 4.3.
Algorithm 1 GRAMMAR-Gen
@@ 4 GRAMMAR
Input: SQL Template Generator $g_{\mathrm{sql}}$, Text Template Generator $g_{\mathrm{t}}$, Semantic criteria for SQL template generation $\mathcal{C}_{\mathrm{sql}}$, Linguistic criteria for text template generation $\mathcal{C}_{\mathrm{t}}$, Database $\mathcal{D}$, Database Schema $\mathcal{S}=\left\{S_{1},S_{2},\ldots,S_{n}\right\}$, where each $S_{i}$ is a schema for a table.
Output: Final evaluation data $Q$
1:  $Q\leftarrow\emptyset$
2:  $\mathcal{S}_{\text{target}}\in\mathcal{S}$
3:  $\{tpl_{\text{sql}}\}\leftarrow g_{\mathrm{sql}}(\mathcal{S}_{\text{target}},\mathcal{C}_{\mathrm{sql}})$
4:  for $tpl_{\mathrm{sql}}\in\{tpl_{\text{sql}}\}$ do
5:     $\{tpl_{\text{t}}\}\leftarrow g_{\mathrm{t}}(tpl_{\mathrm{sql}},\mathcal{C}_{\mathrm{t}})$
6:     $P\leftarrow$ extract placeholders from $tpl_{\mathrm{sql}}$
7:     for each $p\in P$ do
8:        $C\leftarrow$ extract column from $p$
9:        $V_{p}\leftarrow$ query $\mathcal{D}$ for distinct values of $C$
10:     end for
11:     $Comb\leftarrow$ Cartesian product of $\{V_{p}:p\in P\}$
12:     for each $comb\in Comb$ do
13:        $q_{\mathrm{sql}}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{sql}}$
14:        $a\leftarrow$ query $\mathcal{D}(q_{\mathrm{sql}})$
15:        $\{q_{t}\}\leftarrow$ substitute $comb$ into $tpl_{\mathrm{t}}\in\{tpl_{\text{t}}\}$
16:        $Q\leftarrow Q\cup\{(\{q_{t}\},a)\}$
17:     end for
@@ 4.1 Generating Query Templates
The GRAMMAR-Gen process first utilizes database schemas and LLMs to generate SQL query templates and their corresponding textual forms.
@@
Database schema serves as the blueprint that defines how classes of entities are organized and the relations among them. Appendix B demonstrates the reasons why we use database schema for knowledge representation. Let a database schema with a collection of $N$ tables $S=\left\{S_{1},S_{2},\ldots,S_{N}\right\}$, where each $S_{i}$ is a schema for a database table. A table schema $S_{i}$ consists of the table name $T_{i}$, a set of attributes $\left\{A_{1},A_{2},\ldots,A_{M}\right\}$ and a set of constraints, i.e., $S_{i}=\left(T_{i},\left\{A_{1},A_{2},\ldots,A_{m}\right\}\right.,C)$, where the constraint $C$ demonstrates the primary key $PK$ and foreign keys $FK$. Specifically, $FK\left(T_{i}.A_{k}\rightarrow T_{j}.A_{h}\right)$ indicates a foreign key $A_{k}$ in table $T_{i}$ referencing the primary key $A_{h}$ in table $T_{j}$. The schema defines the knowledge structure of an entity, which can be enough to generate query semantics that is not tied to any specific linguistic expressions and sensitive data.
@@
$tpl_{\mathrm{sql}}$ can be instantiated into various text templates, each denoted as $tpl_{t}$. This process, driven by the text template generator $g_{\mathrm{t}}$, leverages linguistic criteria $\mathcal{C}_{\mathrm{t}}$ to transform $tpl_{\mathrm{sql}}$ into natural language forms that align with specified linguistic characteristics, e.g., complexity, length and stylistic nuances  [^4] Examples of $\mathcal{C}_{\mathrm{sql}}$ and $\mathcal{C}_{\mathrm{t}}$ are specified in Appendix C.. The text template generator $g_{\mathrm{t}}$ is also operationalized using a sophisticated language model, such as GPT-4, which is adept at producing diverse linguistic variations of the same semantic content.
$\{tpl_{\text{t}}\}=g_{\text{t}}(tpl_{\text{sql}},\mathcal{C}_{\text{t}}),$
where $\{tpl_{\text{t}}\}$ signifies the resulting list of text templates.